# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ III.4 The quenched-noise approximation for the LIF neuron rate variance
To our knowledge, the rate variance ${v_{\rm R}}$ of a LIF neuron with white+coloured noise input can be analytically computed only in the ${\tau_{\rm c}}\to 0$ limit [12], in the ${\tau_{\rm c}}\to\infty$ limit with no white noise [21], for a perfect integrator [20, 18], for conductance synapses [14], for binary-valued inputs [15] or for oscillating, periodic input [13]. In any case, there are no closed analytical expressions for the problem at hand (see also [22])  [^1] The recent results [49, 53] manage to have an approximated expression for the first-passage time distribution of an active Brownian particle (a non-Markovian random walker) for small $\nu_{\rm W}={\sigma_{\rm c}}^{2}{\tau_{\rm m}}/({\sigma_{\rm I}}^{2}{\tau_{\rm c}})$. Here we derive an expression for the rate variance within the quenched-noise approximation.
@@
We compute ${g_{\rm q}}(|t-t^{\prime}|)$ in the quenched-noise approximation (Appendix G) by using the facts that (i) the effective encoding current $\tilde{\mu}_{t}:={\mu_{\rm I}}+{\sigma_{\rm c}}\xi(t)$ (III.1) obeys an Ornstein-Uhlenbeck process (Eq. (5)) with known stationary distribution $p(\tilde{\mu}_{t})={\cal N}(\tilde{\mu}_{t};{\mu_{\rm I}},{v_{\rm c}})$, and Green function, $p(\tilde{\mu}_{t^{\prime}}|\tilde{\mu}_{t})$; (ii) in the quenched-noise approximation we have that ${r_{\rm q}}{g_{\rm q}}(t-t^{\prime})\simeq\langle\langle{r_{\rm w}}(\tilde{\mu}_{t}){r_{\rm w}}(\tilde{\mu}_{t^{\prime}})\rangle_{p(\tilde{\mu}_{t^{\prime}}|\tilde{\mu}_{t})}\rangle_{p(\tilde{\mu}_{t})}$. Thus, our derivation addresses the impact of pre-synaptic current fluctuations in the post-synaptic temporal correlations, beyond their impact in the post-synaptic mean firing rate. The resulting expression for ${g_{\rm q}}(|t-t^{\prime}|)$ is expected to be valid for resolution timescales $\ell$ satisfying ${\tau_{\rm m}}\ll\ell\ll{\tau_{\rm c}}\ll{T_{\rm c}}$ and for low $\omega:=\exp(-|t-t^{\prime}|/{\tau_{\rm c}})$ and is given by (Appendix G):
${r_{\rm R}}{g_{\rm q}}^{(\text{lo-}\omega)}(t-t^{\prime})$ $\simeq{r_{\rm q}}^{2}\left(1+e^{-|t-t^{\prime}|/{\tau_{\rm c}}}\,Q\right)\qquad\text{for large $|t-t^{\prime}|/{\tau_{\rm c}}$}$ (12)
where $Q$ is a constant that depends on the white LIF properties ${\mu_{\rm I}},{\sigma_{\rm I}},{\tau_{\rm m}}$, and on ${v_{\rm c}}$ (see Eq. (61) in Appendix G).
@@
Figure 6: Temporal distribution function $g(t)$ of readout spikes. Different curves/colors correspond to different values of ${\tau_{\rm c}}$. The points correspond to the numerical estimation from the simulated spike trains (see Sec. C); the colored continuous curves correspond to the quenched approximation ${g_{\rm q}}$ in (12) (please, note that these curves become bordered by a black line for $t>{\tau_{\rm c}}$, in the regime in which they have been derived). Importantly, in order to correct the finite-NIS effect, we have shifted the numerical $g$ points by a quantity, ${\tau_{\rm c}}$-dependent but constant in ${\mu_{\rm I}}$, which is $-\tilde{r_{\rm R}}+{r_{\rm q}}$, where $-\tilde{r_{\rm R}}$ is the numeric estimation of the rate (which is underestimated, due to the finite-NIS effect). To show the effect of such a shift, we report as well the raw, non-shifted numerical estimations (small, void triangles). The gray dashed lines are the hand-waving expression for $g$, (69) in Appendix. H. The dashed and continuous horizontal lines mark, respectively, the values ${v_{\rm q}}+{r_{\rm q}}^{2}$ and ${r_{\rm q}}^{2}$. The black vertical line signals ${\mu_{\rm I}}^{-1}$, while the vertical colored lines mark the corresponding values of ${\tau_{\rm c}}$. It is interesting to note that the hand-waving approximations for $g$ seem to work slightly worst than ${g_{\rm q}}$, as expected. The parameters of the simulation are: ${\mu_{\rm I}}=47.4$, ${\alpha_{\rm V}}=0.9$, ${v_{\rm V}}=9.8$, ${v_{\rm c}}=650.22$, ${\sigma_{\rm I}}=4.618$.
@@
Figure 7: Standard deviation of the firing rate in the unit time interval, $\rm{Std}(\rm{rate})\,{T_{\rm c}}^{1/2}=({v_{\rm R}}{T_{\rm c}})^{1/2}$, vs ${\mu_{\rm I}}$. Different panels correspond to two simulations with different values of ${\tau_{\rm c}}$ (${\tau_{\rm c}}=0.0052$ and $0.1$ for panels A and B respectively). Different curves in each panel correspond to different values of ${\alpha_{\rm V}}=0,0.9$: absence of noise correlations, and highly aligned noise correlations, respectively. The points are the numerical estimations: standard deviations of ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$ across the ${n_{\rm tr}}$ trials (×$\times{T_{\rm c}}^{1/2}$). The error-bars are the confidence interval of the corresponding $\chi^{2}$ variance, with a probability $\alpha_{\chi}=0.9$ (see Appendix C). Continuous lines are the Poisson-quenched hypothesis ${v_{\rm R}}={r_{\rm q}}/{T_{\rm c}}$; the filled, continuous-line boxes are our quenched-noise upper and lower bounds for the variance (13) (upper and lower limits of the box); the void, dashed-line boxes correspond to the upper and lower bounds in the hand-waving approximation (see Appendix H). We observe that in absence of noise correlations when ${\alpha_{\rm V}}=0$, the Poisson approximation is relatively good, specially for low ${\mu_{\rm I}}$. Only large values of ${\sigma_{\rm c}}$ induce positive corrections of the Poisson variance. The quenched-noise approximation captures both negative and positive deviations from Poissonian variance. The rest of the parameters are: ${\tau_{\rm m}}=0.005$, $w=0.450$, ${\sigma_{\rm V}}=1.104$.
@@
We can exploit our expression for ${g_{\rm q}}^{(\text{lo-}\omega)}$ to draw two approximations for the rate variance (see Fig. 6), depending on the value of a temporal cutoff $t_{0}$.
@@
These approximations capture the rate variances in a broad range of ${\mu_{\rm I}}$ and ${\tau_{\rm c}}$ parameters (see Fig. 7 and Appendix F). In particular, the quenched-noise approximation to the variance in Eq. (13) captures both the negative deviations from Poissonian noise $\text{Var}(\text{rate})={r_{\rm R}}/{T_{\rm c}}$ (that originate from the refractory period for $t>t_{0}$, ${\sf C}(t)<0$), and the over-Poissonian deviations (originating from the high-$t$ regime, in which ${\sf C}(t)>0$).
In Appendix H, we develop a further approximation to the firing rate variance, that we will refer to as hand-waving approximation, and which admits a straightforward interpretation of the over-Poissonian enhancement.
@@ IV Noise correlations and SNR enhancement in the encoding-readout model
In this section, we study the behavior of the SNR of the readout (Eq. 4) as a function of the noise correlation stationary variance ${v_{\rm c}}$ (parametrised by ${\alpha_{\rm V}}$) and timescale ${\tau_{\rm c}}$. Crucially, for the comparison between numerical estimations and theoretical approximations for different values of ${\tau_{\rm c}}$, we keep the noise stationary variance ${v_{\rm c}}$ constant and independent on ${\tau_{\rm c}}$, which amounts to consider a ${\tau_{\rm c}}$-dependent count variance ${\sigma_{\rm c}}^{2}=2{\tau_{\rm c}}{v_{\rm c}}$ (see [21]). Taking ${v_{\rm c}}$ constant is the suitable scenario to explore the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit without loss of generality (see Appendix J), which allows us to employ the quenched-noise approximation for the LIF model rate mean and variance, valid in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit.
@@
In IV.1 we first study the behavior of the SNR by comparing the numerical simulations with the following analytical approximations.
For the colored-noise LIF rate ${r_{\rm R}}$ we will use: the two Moreno approximations [17] $r_{\rm M}$ for the firing rate, expected to hold in the limits ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ (Eq. 36), respectively, and for low $\alpha_{\rm M}$ and low ${\mu_{\rm I}}^{2}/({\sigma_{\rm c}}^{2}+{\sigma_{\rm I}}^{2})$; and the quenched-noise approximation ${r_{\rm q}}$ in Eq. (9), expected to hold for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ and for low ${\mu_{\rm I}}^{2}/({\sigma_{\rm c}}^{2}+{\sigma_{\rm I}}^{2})$.
For the colored-noise LIF variance ${v_{\rm R}}$ we will use: (i) The Poisson approximation ${v_{\rm R}}\simeq{r_{\rm R}}/{T_{\rm c}}$, consisting in neglecting the time-correlated part of the temporal correlation ${\sf C}_{-}$ in Eq. (11). This approximation works relatively well in the absence of noise correlations (${\sigma_{\rm c}}=0$), and for low values of ${\mu_{\rm I}}$. Then, the mean rate can be approximated with the Moreno expression ${r_{\rm R}}\simeq r_{\rm M}$ (Poisson-Moreno approximation); and (ii) The quenched-noise approximation for the variance, ${v_{\rm q}}$ in Eq. (9), expected to hold for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ and for low ${\mu_{\rm I}}^{2}/({\sigma_{\rm c}}^{2}+{\sigma_{\rm I}}^{2})$.
IV.1 Signal enhancement, noise enhancement and SNR enhancement
IV.1.1 Signal enhancement

### v2

@@ III.4 The quenched-noise approximation for the LIF neuron rate variance
The rate variance ${v_{\rm R}}$ of a LIF neuron with white+coloured noise input has been analytically computed only in the ${\tau_{\rm c}}\to 0$ limit [12], in the ${\tau_{\rm c}}\to\infty$ limit with no white noise [21], for a perfect integrator [20, 18], for conductance synapses [14], for binary-valued inputs [15] or for oscillating, periodic input [13]. To our knowledge, there are no closed analytical expressions for it (see also [22]) 111[49, 65] obtained approximated expression for the first-passage time distribution of an active Brownian particle (a non-Markovian random walker) for small $\nu_{\rm W}={\sigma_{\rm c}}^{2}{\tau_{\rm m}}/({\sigma_{\rm I}}^{2}{\tau_{\rm c}})$. Here we derive an expression for the rate variance within the quenched-noise approximation.
@@
We compute ${g_{\rm q}}(|t-t^{\prime}|)$ in the quenched-noise approximation (Appendix H) by using the facts that (i) the effective encoding current $\tilde{\mu}_{t}:={\mu_{\rm I}}+{\sigma_{\rm c}}\xi(t)$ obeys an Ornstein-Uhlenbeck process (Eq. (5b)) with known stationary distribution $p(\tilde{\mu}_{t})={\cal N}(\tilde{\mu}_{t};{\mu_{\rm I}},{v_{\rm c}})$ and Green function $p(\tilde{\mu}_{t^{\prime}}|\tilde{\mu}_{t})$; (ii) in the quenched-noise approximation we have that ${r_{\rm q}}{g_{\rm q}}(t-t^{\prime})\simeq\langle\langle{r_{\rm w}}(\tilde{\mu}_{t}){r_{\rm w}}(\tilde{\mu}_{t^{\prime}})\rangle_{p(\tilde{\mu}_{t^{\prime}}|\tilde{\mu}_{t})}\rangle_{p(\tilde{\mu}_{t})}$. Thus, our derivation addresses the impact of pre-synaptic current fluctuations in the post-synaptic temporal correlations, beyond their impact in the post-synaptic mean firing rate. The equation ${r_{\rm q}}{g_{\rm q}}(t-t^{\prime})\simeq\langle\langle{r_{\rm w}}(\tilde{\mu}_{t}){r_{\rm w}}(\tilde{\mu}_{t^{\prime}})\rangle_{p(\tilde{\mu}_{t^{\prime}}|\tilde{\mu}_{t})}\rangle_{p(\tilde{\mu}_{t})}$ is the quenched-noise approximation to ${g_{\rm q}}$, expected to be valid for resolution timescales $\ell$ satisfying ${\tau_{\rm m}}\ll\ell\ll{\tau_{\rm c}}\ll{T_{\rm c}}$. In Appendix H we show that, under the further conditions of small ${v_{\rm c}}$ and low $\omega:=\exp(-|t-t^{\prime}|/{\tau_{\rm c}})$, one can derive an analytic expression for ${g_{\rm q}}$, given by (see Appendix H):
${r_{\rm R}}{g_{\rm q}}^{(\text{lo-}\omega)}(t-t^{\prime})$ $\simeq{r_{\rm q}}^{2}\left(1+e^{-|t-t^{\prime}|/{\tau_{\rm c}}}\,Q\right)\qquad\text{low ${v_{\rm c}}$, large $|t-t^{\prime}|/{\tau_{\rm c}}$}$ (12)
where $Q$ is a constant that depends on the white LIF properties ${\mu_{\rm I}},{\sigma_{\rm I}},{\tau_{\rm m}}$, and on ${v_{\rm c}}$ (see Eq. (99) in Appendix H).
@@
Figure 7: Temporal distribution function $g(t)$ of readout spikes. Different curves/colors correspond to different values of ${\tau_{\rm c}}$. The points correspond to the numerical estimation from the simulated spike trains (see Appendix C); the colored continuous curves correspond to the quenched approximation ${g_{\rm q}}$ in (12) (please, note that these curves become bordered by a black line for $t>{\tau_{\rm c}}$, in the regime in which they have been derived). Importantly, in order to correct the finite-NIS effect, we have shifted the numerical $g$ points by a quantity, ${\tau_{\rm c}}$-dependent but constant in ${\mu_{\rm I}}$, which is $-\tilde{r_{\rm R}}+{r_{\rm q}}$, where $-\tilde{r_{\rm R}}$ is the numeric estimation of the rate (which is underestimated, due to the finite-NIS effect). To show the effect of such a shift, we report as well the raw, non-shifted numerical estimations (small, void triangles). The gray dashed lines are the hand-waving expression for $g$, (107) in Appendix. I. The dashed and continuous horizontal lines mark, respectively, the values ${v_{\rm q}}+{r_{\rm q}}^{2}$ and ${r_{\rm q}}^{2}$. The black vertical line signals ${\mu_{\rm I}}^{-1}$, while the vertical colored lines mark the corresponding values of ${\tau_{\rm c}}$. It is interesting to note that the hand-waving approximations for $g$ seem to work slightly worst than ${g_{\rm q}}$, as expected. The parameters of the simulation are: ${\mu_{\rm I}}=47.4$, $w=0.45$, ${\sigma_{\rm c}}=30.593$, ${\tau_{\rm m}}=0.005$, ${\tau_{\rm c}}=0.7196$.
@@
Figure 8: Standard deviation of the firing rate in the unit time interval, ${\rm Std}({\rm rate})\,{T_{\rm c}}^{1/2}=({v_{\rm R}}{T_{\rm c}})^{1/2}$, vs ${\mu_{\rm I}}$. Different panels correspond to two simulations with different values of ${\tau_{\rm c}}$ (${\tau_{\rm c}}=0.0052$ and $0.1$ for panels A and B respectively). Different curves in each panel correspond to different values of ${\alpha_{\rm V}}=0,0.9$: absence of noise correlations, and highly aligned noise correlations, respectively (see the value of ${\sigma_{\rm c}}$ in the legend). The points are the numerical estimations, and the error-bars are the confidence interval of the corresponding $\chi^{2}$ variance, with a probability $\alpha_{\chi}=0.9$ (see Appendix C). Continuous lines are the Poisson-quenched hypothesis ${v_{\rm R}}={r_{\rm q}}/{T_{\rm c}}$; the filled, continuous-line boxes are our quenched-noise upper and lower bounds for the variance (13) (upper and lower limits of the box); the void, dashed-line boxes correspond to the upper and lower bounds in the hand-waving approximation (see Appendix I). We observe that in absence of noise correlations when ${\alpha_{\rm V}}=0$, the Poisson approximation is relatively good, specially for low ${\mu_{\rm I}}$. Only large values of ${\sigma_{\rm c}}$ induce positive corrections of the Poisson variance. The quenched-noise approximation captures both negative and positive deviations from Poissonian variance. The rest of the parameters are: ${\tau_{\rm m}}=0.005$, $w=0.45$, $N=20$, ${\sigma_{\rm V}}=1.104$.
@@
We exploit our expression for ${g_{\rm q}}^{(\text{lo-}\omega)}$ to draw two approximations for the rate variance, depending on the value of a temporal cutoff $t_{0}$.
@@
These approximations capture the rate variances in a broad range of ${\mu_{\rm I}}$ and ${\tau_{\rm c}}$ parameters (see Fig. 8 and Appendix G). In particular, the quenched-noise approximation to the variance in Eq. (13) captures both the negative deviations from Poissonian noise $\text{Var}(\text{rate})={r_{\rm R}}/{T_{\rm c}}$ (that originate from the refractory period for $t<t_{0}$, ${\sf C}(t)<0$), and the over-Poissonian deviations (originating from the high-$t$ regime, in which ${\sf C}(t)>0$).
Importantly, in Appendix I, we develop a further approximation to the firing rate variance, that we will refer to as hand-waving approximation, and which admits a straightforward interpretation of the over-Poissonian enhancement. In the hand-waving approximation approximation, the expressions for the temporal distribution function and for the rate variance are as in the above equations, for a different value of the constant $Q$, $Q^{({\rm hand})}:={u_{\rm q}}/{r_{\rm q}}$, where ${u_{\rm q}}$ is, in analogy with Eq. (10), the variance of the white-noise LIF activation function over the stationary distribution of ${\mu_{\rm I}}$’s, ${u_{\rm q}}:=\left<{r_{\rm w}}^{2}(\tilde{\mu})\right>_{{\cal N}(\tilde{\mu};\mu_{I},{v_{\rm c}})}-{r_{\rm q}}^{2}$. Despite the simplicity of its demonstration, the hand-waving approximation for the rate variance is almost as accurate as its more accurate variant derived in Appendix H (Fig. 7). As in the previous subsection, the above analytic expressions for ${g_{\rm q}},{v_{\rm q}}$ require, besides the validity condition of the quenched-noise approximation ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, the further condition of small ${v_{\rm c}}$ (see details in H).
@@ IV Noise correlations and SNR enhancement in the encoding-readout model
In this section, we study the behavior of the SNR, Eq. (4) as a function of the noise correlation stationary variance ${v_{\rm c}}$ (parametrised by ${\alpha_{\rm V}}$) and timescale ${\tau_{\rm c}}$. Crucially, for the comparison between numerical estimations and theoretical approximations for different values of ${\tau_{\rm c}}$, we keep the noise stationary variance ${v_{\rm c}}$ constant (where, throughout this section, by constant we mean independent of ${\tau_{\rm c}}$), which amounts to consider a ${\tau_{\rm c}}$-dependent count variance ${\sigma_{\rm c}}^{2}=2{\tau_{\rm c}}{v_{\rm c}}$ (see [21]). Taking ${v_{\rm c}}$ constant is the suitable scenario to explore the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit without loss of generality (see Appendix L).
The fact that ${\sigma_{\rm c}}$ depends on ${\tau_{\rm c}}$ for constant ${v_{\rm c}}$ is an arbitrary choice of the correlated noise amplitude that we probe in our simulations for each value of ${\tau_{\rm c}}$. Since ${\tau_{\rm c}},{\sigma_{\rm c}}$ do not vary independently (in the so called first simulation set analyzed in this section, see Appendix C), we decided to set the spike count variance ${\sigma_{\rm c}}$ to depend on the independent variable ${\tau_{\rm c}}$ as ${\sigma_{\rm c}}=(2{\tau_{\rm c}}{v_{\rm c}})^{1/2}$ (being ${v_{\rm c}}$ a ${\tau_{\rm c}}$-independent proportionality constant), rather than taking the alternative option (i.e., ${v_{\rm c}}$ depending on ${\tau_{\rm c}}$ as ${v_{\rm c}}={\sigma_{\rm c}}^{2}/2{\tau_{\rm c}}$, being ${\sigma_{\rm c}}$ a ${\tau_{\rm c}}$-independent proportionality constant) because this scenario makes more sense when studying the large-${\tau_{\rm c}}$ limit (see Appendix L).
We provide an alternative analysis of the SNR in a complementary probing variables scheme, in which we fix ${\tau_{\rm c}}$ and vary ${\sigma_{\rm c}}$ (instead of fixing ${v_{\rm c}}$ and varying ${\tau_{\rm c}}$), in Appendix F.
IV.1 Use of the analytical approximations
@@
Below in subsection IV.2, we present a study of the behavior of the SNR as a function of ${\tau_{\rm c}}$ by comparing the numerical simulations of ${r_{\rm R}}^{(s)}$, ${v_{\rm R}}^{(s)}$ and of the SNR, with the following analytical approximations.
In the region ${\tau_{\rm c}}\ll{\tau_{\rm m}}$, the only analytical solution that we need in order to explain the numerical estimations of the SNR, is the Moreno et al solution [17] for the rate, $r_{\rm M}$, in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ limit (Eq. (10) of [17]). Two different conditions make possible to approximate the ${r_{\rm R}}$, ${v_{\rm R}}$ and SNR, disposing only of this analytical expression. First, because in the limit ${\tau_{\rm c}}\ll{\tau_{\rm m}}$, and for fixed ${v_{\rm c}}$, ${\sigma_{\rm c}}$ is small enough for our simulations to be in the validity regime of the Moreno solution (low $\alpha_{\rm M}={\sigma_{\rm c}}^{2}/{\sigma_{\rm I}}^{2}$) and, hence, the Moreno et al solution for ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ reproduces correctly the numerical simulations for ${r_{\rm R}}$ (see Fig. 9-A). Second, because, for low enough ${\tau_{\rm c}}$, hence low enough ${\sigma_{\rm c}}$, we can neglect the over-Poissonian corrections to the variance ${v_{\rm R}}$ in Eq. (11) (see the ${\alpha_{\rm V}}=0$ case in Fig. 8). In sum, we simply take ${v_{\rm R}}{T_{\rm c}}\simeq{r_{\rm R}}\simeq r_{\rm M}$, where $r_{\rm M}$ is the Moreno et al approximation (we will refer to this approximation for the readout rate variance as the Poisson-Moreno approximation). We construct our analytical expression for the SNR with these expressions for ${r_{\rm R}}$ and ${v_{\rm R}}$.
For ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, none of these conditions hold any more. Although in principle we dispose of the alternative Moreno et al solution for $r_{\rm M}$, valid for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ (that we report in Eq. 41), it is no longer possible to exploit this solution to explain the behavior of the SNR for large ${\tau_{\rm c}}$ and constant ${v_{\rm c}}$. Indeed, for large enough ${\tau_{\rm c}}$, hence ${\sigma_{\rm c}}$, the simulation parameters lie outside the validity limit (low $\alpha_{\rm M}$) of the Moreno solution and, consequently, $r_{\rm M}$ differs from the numerical estimation $\tilde{r_{\rm R}}$. Second, and more importantly, the over-Poissonian corrections of the rate variance cannot be neglected any longer. Therefore, we cannot approximate ${r_{\rm R}}$ with $r_{\rm M}$, and we cannot use any longer the relation ${v_{\rm R}}{T_{\rm c}}\simeq{r_{\rm R}}\simeq r_{\rm M}$. The quenched-noise approximation for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ provides us with a solution for both problems. First, ${r_{\rm q}}$ captures well the behavior of ${r_{\rm R}}$ also in this regime. Second, the quenched-noise approximation provides us with an equation for the rate variance as well (Eq. (13), or the simpler hand-waving quenched-noise approximations Eqs. (109,110)). We illustrate these facts in the next subsection, and particularly in Fig. 9, in which we compare the Poisson-Moreno approximation with the quenched-noise approximation.
IV.2 Signal enhancement, noise enhancement and SNR enhancement
IV.2.1 Signal enhancement